AnimeGhoul8863:

The table below shows four systems of equations: System 1 System 2 System 3 System 4 4x − 5y = 2 3x − y = 8 4x − 5y = 2 3x − 2y = 1 4x − 5y = 2 3x − 8y = 4 4x − 5y = 2 10x − 9y = 4 Which pair of systems will have the same solution? System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to two times the second equation in system 1 System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to three times the second equation in system 1 System 2 and system 4, because the second equation in system 4 is obtained by adding the first equation in system 2 to two times the second equation in system 2 System 2 and system 4, because the second equation in system 4 is obtained by adding the first equation in system 2 to three times the second equation in system 2

AnimeGhoul8863:

4x − 5y = 2= system1 3x − y = 8=system 1 4x − 5y = 2= System 2 3x − 2y = 1 =System 2 4x − 5y = 2= System 3 3x − 8y = 4= System 3 4x − 5y = 2= System4 10x − 9y = 4= System 4

Ultrilliam:

@Shadow @Vocaloid

Ultrilliam:

@dude

Ultrilliam:

@563blackghost Can you help this person when you get the chance?

563blackghost:

Hello! Let me take a look at your question :)

AnimeGhoul8863:

OK thank you so much

AnimeGhoul8863:

if a pic would make more sense give me 2 mins to get one

AnimeGhoul8863:

1 attachment
563blackghost:

No I understand it just trying to figure it out. It's got multiple steps. The best thing is to try out each answer choice. So I am doing that.

AnimeGhoul8863:

ok thanks for all the effort ur putting in

563blackghost:

Let's try the first one. ` System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to two times the second equation in system 1` So the second equation in system 3 is \(\bf{3x-8y=4}\). It says it is obtained by adding the first equation of system 1 (which is \(\bf{4x-5y=2}\) to 2 times the second equation of 1 (which is \(\bf{3x-y=8}\) so we try this out. Apply 2 times. \(\large\bf{3x-y=8 \rightarrow 6x-2y=16}\) Now add this to the first equation. \(\large\bf{10x-7y=18 }\) This is incorrect. Let's try the next one.

563blackghost:

Second goes by... `System 1 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 1 to three times the second equation in system 1` Second equation of system 3: \(\bf{3x-8y=4}\) First equation in system 1: \(\bf{4x-5y=2}\) 3 times the second equation in 1: \(\bf{3x-2y=1 \rightarrow 9x - 6y = 3}\) Now add these. \(\large\bf{13x-11y=5}\) Is this correct or incorrect?

563blackghost:

Does `13x-11y=5` equal to `3x-8y=4`?

563blackghost:

u there @AnimeGhoul8863 ?

AnimeGhoul8863:

im sorry my mom was talking to me for a moment its my grandparents bday so i had to call them but im back now

563blackghost:

okie that's fine :) The question I asked is with the explanation of the second answer (which is above) is the total equal to the seond equation of system 3 (\(\bf{3x-8y=4}\))?

AnimeGhoul8863:

My mind says yes but my gut says no

563blackghost:

correct they are not equal so B is wrong.

563blackghost:

Now on to the next XD

AnimeGhoul8863:

im thinking the answer of this question is c but i still cant tell

563blackghost:

The third choice tells us... ` System 2 and system 4, because the second equation in system 4 is obtained by adding the first equation in system 2 to two times the second equation in system 2` Second equation of system 4 is \(\bf{10x-9y=4}\) It tells us that the first equation in system 2 (which is \(\bf{4x-5y=2}\)) is added to twice the second equation in system 2 (which is \(\bf{3x-2y=1}\)). We follow the process. Twice the second equation in system 2 \(\large\bf{3x-2y=1 ~~\rightarrow ~~ 6x-4y=2}\) Added it to the first equation of system 2 \(\bf{4x-5y=2}\) . \(\large\bf{10x-9y=4}\) Does the second equation of system 4 \(\bf{10x-9y=4}\) equal to \(\bf{10x-9y=4}\)?

AnimeGhoul8863:

yes

AnimeGhoul8863:

so the answer is C

563blackghost:

Correct :D

AnimeGhoul8863:

YEAH FINALLY GOT THE ANSWER THANK YOU SOOOOOOOOOOOOOOOO MUCH!!!!!!!!!!!!!!!!!!!!!!!!!

563blackghost:

Your welcome ^.^